Problem 5-4 Calculating Interest Rates [L03] Solve for the unknown interest rate in each of the following (Do not round intermediate calculations and round your final answers to 2 decimal places. (e.g., 32.16)): Present Value Years Interest Rate Future Value $ 232 $ 200 % 320 17 847 35,000 8 34,26120 143,410 261,665
Answer #1

i = [future value /present value ]^1/n    -1

a)i = [232/200]^1/3   -1

     =1.0507-1

     = .0507 or 5.07%

b)i =[847/320]^1/17 - 1

       = [2.64688]^.05882 - 1

       = 1.0589-1

         = .0589 or 5.89%

c) i= [143410/35000]^1/18 -1

   =1.0815-1

   = .0815 or 8.15%

d)i = [261665/34261]^1/20   -1

    = [7.63740]^.05     -1

     = 1.1070-1

     = .1070 or 10.70%